OpenStudy (anonymous):

Write the equation of a circle with a center at (–4, 2) and a radius of 2.2.

OpenStudy (anonymous):

(x – 4)2 + (y + 2)2 = 4.84 (x + 4)2 + (y – 2)2 = 4.84 (x + 4)2 + (y – 2)2 = 4.4 (x + 4)2 + (y – 2)2 = 2.2

OpenStudy (zzr0ck3r):

the equation of a circle with center \((a,b)\) and radius \(r\) is \((x-a)^2+(y-b)^2=r^2\)